12. The artificial sweetener NutraSweet is the compound aspartamine which has the molecular formula C14H18N,Os. a) What is
Answer #1

atomic weight of C =12 g, H=1 gm, O=16 gm, N =14 gm

molecular formula = C14H18N2O5

molecular mass M = sum of each mass of atom present in the molecule

molecular mass of aspartamine =mass of C +mass of Hydrogen +mass of O +mass of N

MC14H18N2O5= 14*12 g+18*1 g+14*2 g+16*5 g

MC14H18N2O5= 168 g+18g +28g+80g

MC14H18N2O5= 294 g

1 mole of any substance is equal to its gram molecular mass and contain 6.023 *1023 molecules

a) mass of 1 mol of aspartamine =294 g

b) mass taken 6.22g

moles n = mass taken(m)/molecular mass (M)

n= m/M

n= 6.22 g/294g= 0.0211 moles

c) 1 mole of C14H18N2O5 weighs 294 g

0.245  mol of aspartamine weighs = 294 g*0.245 = 72.03 g

d)1mole or 294 g aspartamine  contain 6.023 *1023 molecules

1gm of aspartamine  contain 6.023 *1023/294 molecules

4.28mg = 4.28*10-3 g  of aspartamine  contain =4.28*10-3 g *6.023 *1023/294 molecules [1gm =1000mg]

4.28*10-3 g  of aspartamine  contain =8.769 *1018molecules